# Evaluating language models as risk scores

### v1

@@ Abstract
Current question-answering benchmarks predominantly focus on accuracy in realizable prediction tasks. Conditioned on a question and answer-key, does the most likely token match the ground truth? Such benchmarks necessarily fail to evaluate language models’ ability to quantify outcome uncertainty. In this work, we focus on the use of language models as risk scores for unrealizable prediction tasks. We introduce folktexts, a software package to systematically generate risk scores using large language models, and evaluate them against benchmark prediction tasks. Specifically, the package derives natural language tasks from US Census data products, inspired by popular tabular data benchmarks. A flexible API allows for any task to be constructed out of 28 census features whose values are mapped to prompt-completion pairs. We demonstrate the utility of folktexts through a sweep of empirical insights on 16 recent large language models, inspecting risk scores, calibration curves, and diverse evaluation metrics. We find that zero-shot risk sores have high predictive signal while being widely miscalibrated: base models overestimate outcome uncertainty, while instruction-tuned models underestimate uncertainty and generate over-confident risk scores.
@@ 1 Introduction
The evaluation of LLMs to date, however, has predominantly focused on accuracy metrics, often in realizable tasks where there is a unique correct label for each data point. Such benchmarks necessarily cannot speak to the use of a language model as risk scores. A model can have high utility in well-defined question-answering tasks while being wildly miscalibrated. Indeed, growing empirical evidence suggests that instruction-tuned language models tend to be over-confident in the probability that they assign to their preferred answer [9, 10].
@@ 1.1 Our contributions
We contribute an open-source software package, called folktexts, [^1] Package and results available at: https://github.com/socialfoundations/folktexts that provides datasets and tools to evaluate statistical properties of LLMs on classification benchmarks. We show-case these functionalities with a sweep of new empirical insights on the risk scores produced by 16 recently proposed LLMs.
Classical metrics and tools for evaluating statistical properties of machine learning models rely on a standard type signature of numeric features $X$, labels $Y$, and risk scores $R$. As language models interface with the world strictly through natural text, they are incompatible with a rich repertoire of open-source libraries, benchmarks and evaluation tools. The folktexts package is designed to bridge this gap and offer a systematic way to translate between the natural language interface and the standard $(X,Y,R)$ statistical machine learning type signature. Figure 1 illustrates the workflow for producing risk scores using LLMs.
Inspired by the popular folktables package [11], folktexts derives natural-text machine learning datasets from US Census data products, specifically, the American Community Survey (ACS) [12]. Folktexts systematically constructs prompts from the survey responses of more than 3.2 million individuals, using the US Census codebook, and the ACS questionnaire as a reference. Risk scores are extracted from the model’s output token probabilities using a standard question-answering interface. The package offers five pre-specified question-answering benchmark tasks that are ready-to-use with any language model. These tasks can be customized through a flexible API where any census data attribute can be used either as a feature or as the target variable in a prediction task (see Table A5 for the initial set of 28 mapped attributes). Furthermore, evaluations can easily be performed over subgroups of the population to conduct algorithmic fairness audits.
@@
We contribute a sweep of empirical findings based on our package. We evaluate 16 recently proposed LLMs, with sizes ranging from 2B parameters to 70B parameters. Our study demonstrates how inspecting risk scores of LLMS on underspecified prediction tasks reveals new insights that can not be deduced from inspecting accuracy alone. The main findings are summarized as follows:
  • Models’ output token probabilities have strong predictive signal, [^2] We measure predictive signal using the area under the receiver operating characteristic curve (AUC). but they are widely miscalibrated.
  • The failure modes of models are different: Base models consistently overestimate uncertainty and produce under-confident risk scores, while their instruction-tuned counterparts underestimate uncertainty and produce over-confident risk scores.
  • As a result instruction-tuning worsens calibration error on tasks with high outcome uncertainty, despite generally improving predictive signal and accuracy.
[…]
@@
In Section 2 we provide necessary background on risk scores and calibration in statistical machine learning. In Section 3 we extend this background to the application to language models, providing various design choices around prompting templates and ways to extract risk scores from language models. In Section 4 we evaluate 16 recent LLMs on 5 proposed benchmark tasks and summarize empirical findings.
@@ 1.3 Related work
The use of LLMs for decision-making has seen increasing interest as of late. Hegselmann et al. [20] show that a Bigscience T0 11B model [21] surpasses the predictive performance (AUC) of supervised learning baselines in the very-few-samples regime. The authors find that fine-tuning an LLM outperforms fitting a standard statistical model on a variety of small to medium-size tasks (from 768 to 48K samples) up to training set sizes in the 10s to 100s of samples.
Related work by Tamkin et al. [1] uses a similar tabular data set-up to study potential biases of the closed-source Claude 2.0 model. The authors generate hypothetical individual information for a variety of decision scenarios, and find that sensitive demographic information (such as age, race and gender) is actively used by the model for decision-making. Some attributes affect the decision positively (e.g., higher chance of approving a small business loan for minorities) while others affect it negatively (e.g., lower chance for older aged individuals).
It is unclear to which degree using synthetic or unrealistic data can affect the studies’ validity. Our proposed benchmarks enable similar studies to be conducted using real-world Census data. Moreover, related literature generally focuses on studying accurate class predictions by specific language models. We study over a dozen state-of-the-art language models, and discern which trends are actually consistent across all models, and which are constrained to a single model or model family. Further, we focus on a particularly neglected research avenue: analyzing LLM risk scores instead of discrete token answers, and whether they are accurate and calibrated.
@@
Calibration is a widely studied concept in the literature on forecasting in statistics and econometrics with a venerable history [22, 23, 24, 25, 26, 27, 28]. Recent years have seen a surge in interest in calibration in the context of deep learning [29]. Calibration of LLMs has been studied on diverse question-answering benchmarks, ranging from sentiment classification, knowledge testing, and mathematical reasoning to multi-task benchmarks [e.g., 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. What differentiates our work is that we study calibration in naturally underspecified prediction tasks. This requires even the most accurate models to accurately reflect non-trivial probabilities over outcomes to be calibrated.
@@
Beyond task-calibration, calibration at the word and token-level has been explored in the context of language generation [e.g., 48]. Others have focused on connections to hallucination [49], and expressing uncertainty in natural language [50, 51, 52]. Related to our work, Lin et al. [53] emphasize the inherent outcome uncertainty in language generation. Focusing on generation at the word or token level poses the challenge of measuring a high-dimensional probability distribution. Considering binary classification tasks has the practical advantage of circumventing this problem.
@@ 2 Preliminaries
We will focus on binary prediction throughout this work. Let $Y\in\{0,1\}$ be a random variable indicating the outcome of an event the learner wishes to predict. We use the shorthand notation $f_{\theta}(x)$ to denote the model’s estimate of the probability that $Y=1$, given context $X=x$. Following standard terminology we will refer to $f_{\theta}:\mathcal{X}\rightarrow[0,1]$ as score function, and its output $f_{\theta}(x)$ as risk score. There are various dimensions along which to evaluate risk scores.
Calibration.
@@
where data points are grouped into $M$ equally spaced bins $B_{m}$ according to their score values. We use $M=10$ in our evaluation, which is a commonly used value [57]. We also provide a quantile-based measure of ECE, as well as Brier score [22] to allow for a more complete picture. Furthermore, we use reliability diagrams [24, 28] to aid a visual interpretation of calibration. These diagrams plot expected sample accuracy as a function of confidence. Any deviation from a perfect diagonal represents miscalibration.
@@
Predictive signal.
From risk scores to accuracy.
@@
The classifier $c$ that minimizes the misclassification error $\operatorname{\mathbb{E}}\mathds{1}\{c(X)\neq Y\}=\mathbb{P}\{c(X)\neq Y\}$ is given by $c^{*}(x)=\mathds{1}\{f^{*}(x)>0.5\}$, where $f^{*}$ is the Bayes optimal scoring function. In the following when we use accuracy we refer to the fraction of correct predictions after thresholding. If not specified otherwise we use $\tau=0.5$. This corresponds to an argmax operator applied to the class probabilities. It is important to note that a classifier can achieve perfect accuracy even when derived from a suboptimal scoring function. Thus, accuracy alone provides an incomplete picture of a model’s ability to express uncertainty.
3 Evaluating lanaguge models as risk scores
We are interested in the performance of LLMs as predictors in unrealizable binary prediction tasks. We prompt models with feature values $x$ to elicit risk scores $r$ and then evaluate these scores against ground truth labels $y$. Recall the illustration of the work flow in Figure 1.
@@ 3.1 Prediction tasks
We construct natural language prediction tasks from the American Community Survey (ACS) Public Use Microdata Sample (PUMS). [^3] https://www.census.gov/programs-surveys/acs/microdata.html The data contains survey responses of about 3.2 million anonymous individuals, carefully curated to offer statistical insights into the population of the United States. We refer to the data as Census data. The Census data contains demographic attributes, as well as information related to income, employment, health, transportation, and housing. Prediction tasks are defined by selecting a subset of attributes to define the features and one attribute to be the label. We threshold continuous target variables and bin multi-class predictions to obtain a binary classification task. To enable straightforward comparison with existing tabular benchmarks, we consider natural text analogues to the tasks in the popular folktables benchmark package [11]. Appendix B describes each task in further detail.
@@
Prediction tasks on human populations typically come with natural outcome uncertainty, meaning that the target label is not uniquely determined by the input features. The fewer features are provided the higher the uncertainty in the outcome. We take this to our advantage to systematically evaluate model calibration. Namely, to be calibrated, a risk score has to reflect both model uncertainty and uncertainty inherent to the prediction task. In fact, the optimal predictor would often output low-confidence answers. In contrast, prevailing question-answering benchmarks always require high confidence for an accurate model to be calibrated. Underspecified, non-realizable prediction tasks allow us to circumvent such potential confounding in calibration evaluations.
@@ 3.2 Extracting risk scores
To extract risk scores from LLMs we map each inference problem to a natural text prompt. For a given datapoint, we specify the classification task in the prompt and extract class probabilities from the model’s next token probabilities, similar to traditional question-answering interfaces. The prompt consists of three components:
@@
    ” This step is typically not needed in supervised classification tasks, because the training data implicitly defines the population. However, for LLMs this is particularly important for evaluating calibration outside the realizable setting, as risk score can not in general simultaneously be calibrated to different populations. Skipping this step could provide insights related to alignment [42, 43] rather than calibration.
@@
Using the constructed prompts, we query the models and extract scores from the next token probabilities for the choice labels A,B as $r_{i}=\mathbb{P}(\texttt{A})/(\mathbb{P}(\texttt{A})+\mathbb{P}(\texttt{B}))$, following the methodology of standard question-answering benchmarks [36, 42]. As LLMs are known to have ordering biases in multiple-choice question-answering [45, 62], we evaluate responses on all choice orderings and average the resulting scores. We refer to Xiong et al. [33] for an overview on alternative design choices on how to elicit confidence scores from LLMs.
@@ 3.3 The folktexts package
The folktexts package is designed to offer a flexible interface between tabular prediction tasks and natural language question-answering tasks in order to extract risk scores from LLMs. The package makes available the ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, ACSTravelTime prediction tasks [11] as natural language benchmarks, together with various functionalities to customize the task definitions and subsample the reference data. The set of attributes available to define the features and label can be found in the ACS PUMS data dictionary.444 https://www.census.gov/programs-surveys/acs/microdata/documentation.html
In addition to providing a reproducible way to extract risk scores from LLMs, folktexts also offers pre-implemented evaluation metrics to benchmark and compare the calibration and accuracy of LLMs. To offer researchers a quick overview over model performance the package also offers functionalities to plot calibration curves for visual inspection of results. The package is easy to use within a python notebook, or from the command line. Further details on the design choices and architecture are available in Appendix C. The next section demonstrates the utility of the package.

### v3

@@ Abstract
Current question-answering benchmarks predominantly focus on accuracy in realizable prediction tasks. Conditioned on a question and answer-key, does the most likely token match the ground truth? Such benchmarks necessarily fail to evaluate LLMs’ ability to quantify ground-truth outcome uncertainty. In this work, we focus on the use of LLMs as risk scores for unrealizable prediction tasks. We introduce folktexts, a software package to systematically generate risk scores using LLMs, and evaluate them against US Census data products. A flexible API enables the use of different prompting schemes, local or web-hosted models, and diverse census columns that can be used to compose custom prediction tasks. We evaluate 17 recent LLMs across five proposed benchmark tasks. We find that zero-shot risk scores produced by multiple-choice question-answering have high predictive signal but are widely miscalibrated. Base models consistently overestimate outcome uncertainty, while instruction-tuned models underestimate uncertainty and produce over-confident risk scores. In fact, instruction-tuning polarizes answer distribution regardless of true underlying data uncertainty. This reveals a general inability of instruction-tuned LLMs to express data uncertainty using multiple-choice answers. A separate experiment using verbalized chat-style risk queries yields substantially improved calibration across instruction-tuned models. These differences in ability to quantify data uncertainty cannot be revealed in realizable settings, and highlight a blind-spot in the current evaluation ecosystem that folktexts covers.
@@ 1 Introduction
The evaluation of LLMs to date, however, has predominantly focused on accuracy metrics, often in realizable tasks where there is a unique correct label for each data point. Such benchmarks necessarily cannot speak to the use of language models as risk score estimators. A model can have high utility in well-defined question-answering tasks while being wildly miscalibrated. In fact, while accuracy corresponds to knowledge over the expected answer, proper uncertainty quantification corresponds to knowledge over the variance over answers.
@@ 1.1 Our contributions
We contribute an open-source software package, called folktexts, [^1] Package and results available at: https://github.com/socialfoundations/folktexts that provides datasets and tools to evaluate statistical properties of LLMs as risk scorers. We show-case its functionalities with a sweep of new empirical insights on the risk scores produced by 17 recently proposed LLMs.
The folktexts package offers a systematic way to translate between the natural language interface and the standard machine learning type signature. It translates prediction tasks defined by numeric features $X$ and labels $Y$ into natural text prompts and extracts risk scores $R$ from LLMs. This opens up a rich repertoire of open-source libraries, benchmarks and evaluation tools to study their statistical properties. Figure 1 illustrates the workflow for producing risk scores using LLMs.
For benchmarking risk scores, we need ground truth samples from a known probability distribution. Inspired by the popular folktables package [9], folktexts builds on US Census data products, specifically, the American Community Survey (ACS) [10], collecting information about more than 3.2 million individuals representative of the US population. Folktexts systematically constructs prediction tasks and prompts from the individual survey responses using the US Census codebook, and the ACS questionnaire as a reference. Risk scores are extracted from the language model’s output token probabilities using a standard question-answering interface. The package offers five pre-specified question-answering benchmark tasks that are ready-to-use with any language model. A flexible API allows for a variety of natural language tasks to be constructed out of 28 census features whose values are mapped to prompt-completion pairs (features detailed in Table A5). Furthermore, evaluations can easily be performed over subgroups of the population to conduct algorithmic fairness audits.
@@
We contribute a sweep of empirical findings based on our package. We evaluate 17 recently proposed LLMs, with sizes ranging from 2B parameters to 141B parameters. Our study demonstrates how inspecting risk scores of LLMs on underspecified prediction tasks reveals new insights that cannot be deduced from inspecting accuracy alone. The main findings are summarized as follows:
  • Models’ output token probabilities have strong predictive signal [^2] We measure predictive signal using the area under the receiver operating characteristic curve (AUC) [11]. but are widely miscalibrated.
  • The failure modes of models are different: Multiple-choice answer probabilities generated by base models consistently overestimate outcome uncertainty, while instruction-tuned models underestimate uncertainty and produce over-confident risk scores.
  • Using a verbalized chat-style prompt results in materially different answer distributions, with significantly improved calibration for instruction-tuned models, accompanied by a small but consistent decrease in predictive power.
  • Instruction-tuning generally worsens calibration of multiple-choice answers, but improves calibration of verbalized numeric answers.
@@
In Section 2 we provide necessary background on risk scores and calibration in statistical machine learning. In Section 3 we extend this background to the application to language models, providing various design choices around prompting templates and ways to extract risk scores from language models. In Section 4 we evaluate 17 recent LLMs on 5 proposed benchmark tasks and summarize empirical findings.
@@ 1.3 Related work
The use of LLMs for decision-making has seen increasing interest as of late. Hegselmann et al. [19] show that a Bigscience T0 11B model [20] surpasses the predictive performance (AUC) of supervised learning baselines in the very-few-samples regime. The authors find that fine-tuning an LLM outperforms fitting a standard statistical model on a variety of tasks up to training set sizes in the 10s to 100s of samples. Related work by Slack and Singh [21] shows that providing task-specific expert knowledge via instructions in context can lead to significant improvements in model predictive power. Tamkin et al. [1] generate hypothetical individual information for a variety of decision scenarios, and analyze how language models’ outputs change when provided different demographic data. Some attributes are found to positively affect the model’s decision (e.g., higher chance of approving a small business loan for minorities) while others affect it negatively (e.g., lower chance for older aged individuals).
A separate research thread considers how to leverage LLMs to model human population statistics. Argyle et al. [22] evaluate whether GPT-3 can faithfully reproduce political party preferences for different US subpopulations, and conclude that model outputs can accurately reflect a variety of correlations between demographics and political preferences. Other works use a similar methodology to model the distribution of human opinions on different domains [23, 24, 25]. Aher et al. [26], Dillion et al. [27] use LLMs to reproduce popular psychology experiments on human moral judgments, and confirm there’s good alignment between human answers and LLM outputs. Literature on modelling human population statistics generally focuses on using LLMs to obtain accurate survey completions. That is, given demographic information on an individual, what was their response to a specific survey question? This methodology often ignores the fact that individuals described by the same set of demographic features will realistically give different answers. An accurate model would not only provide the highest likelihood answer, but also a measure of uncertainty corresponding to the expected variability within a sub-population. Our work tackles this arguably neglected research avenue: Analyzing LLM risk scores instead of discrete token answers, and whether they are accurate and calibrated to human populations.
@@
Calibration is a widely studied concept in the literature on forecasting in statistics and econometrics with a venerable history [28, 29, 30, 31, 32]. Recent years have seen a surge in interest in calibration in the context of deep learning [33]. Calibration of LLMs has been studied on diverse question-answering benchmarks, ranging from sentiment classification, knowledge testing, and mathematical reasoning to multi-task benchmarks [e.g., 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. What differentiates our work is that we study calibration in naturally underspecified prediction tasks. This requires even the most accurate models to accurately reflect non-trivial probabilities over outcomes to be calibrated.
@@
Beyond task-calibration, calibration at the word and token-level has been explored in the context of language generation [e.g., 50]. Others have focused on connections to hallucination [51], and expressing uncertainty in natural language [52, 53, 54]. Related to our work, Lin et al. [55] emphasize the inherent outcome uncertainty in language generation. Focusing on generation at the word or token level poses the challenge of measuring a high-dimensional probability distribution. Considering binary classification tasks has the practical advantage of circumventing this problem.
@@ 2 Preliminaries
We will focus on binary prediction throughout this work. Let $Y\in\{0,1\}$ be a random variable indicating the outcome of an event the learner wishes to predict. We use the shorthand notation $f_{\theta}(x)$ to denote the model’s estimate of the probability that $Y=1$, given context $X=x$. Following standard terminology we will refer to $f_{\theta}:\mathcal{X}\rightarrow[0,1]$ as score function, and its output $f_{\theta}(x)$ as risk score. There are various dimensions along which to evaluate risk scores by comparing them against samples of the reference population $\mathcal{P}$.
2.1 Calibration
@@
where data points are grouped into $M$ equally spaced bins $B_{m}$ according to their score values. We use $M=10$ in our evaluation, which is a commonly used value [59]. We also provide a measure of ECE over quantile-based score bins, as well as Brier score [28] to allow for a more complete picture. Furthermore, we use reliability diagrams [30] to aid a visual interpretation of calibration. These diagrams plot expected sample accuracy as a function of confidence. Any deviation from a perfect diagonal represents miscalibration.
@@
2.2 Predictive performance
@@
The classifier $c$ that minimizes the misclassification error $\operatorname{\mathbb{E}}\mathds{1}\{c(X)\neq Y\}=\mathbb{P}\{c(X)\neq Y\}$ is given by $c^{*}(x)=\mathds{1}\{f^{*}(x)>0.5\}$, where $f^{*}$ is the Bayes optimal scoring function. In the following, when we use accuracy we refer to the fraction of correct predictions after thresholding. If not specified otherwise we use $\tau=0.5$. This corresponds to an argmax operator applied to the class probabilities. It is important to note that a classifier can achieve perfect accuracy even when derived from a suboptimal scoring function. Thus, accuracy alone provides an incomplete picture of a model’s ability to express uncertainty.
Instance ranking.
3 Evaluating language models as risk scores
We are interested in the ability of LLMs to express natural uncertainty in outcomes. Therefore, we construct unrealizable binary prediction tasks, and test the model’s ability to reflect natural variations in underspecified individual outcomes. Specifically, we prompt models with feature values $x$ to elicit risk scores $r$ and then evaluate these scores against ground truth labels $y$ (see Figure 1).
@@ 3.1 Prediction tasks
We construct natural language prediction tasks from the American Community Survey (ACS) Public Use Microdata Sample (PUMS). [^3] https://www.census.gov/programs-surveys/acs/microdata.html The data contains survey responses of about 3.2 million anonymous individuals, carefully curated to offer statistical insights into the population of the United States. We refer to the data as Census data. The Census data contains demographic attributes, as well as information related to income, employment, health, transportation, and housing. Prediction tasks are defined by selecting a subset of attributes to define the features and one attribute to be the label. We threshold continuous target variables and bin multi-class predictions to obtain a binary classification task. Specifically, we test models on their ability to reflect natural variations in the outcome across the benchmark population. To enable straightforward comparison with existing tabular benchmarks, we consider natural text analogues to the tasks in the popular folktables benchmark package [9]. Appendix B describes each task in further detail.
@@
Prediction tasks on human populations typically come with natural outcome uncertainty, meaning that the target label is not uniquely determined by the input features (also known as aleatoric uncertainty [60]). The fewer features are provided the higher the uncertainty in the outcome. We take this to our advantage to systematically evaluate calibration. Namely, to be calibrated, a risk score has to reflect both model uncertainty and uncertainty inherent to the prediction task. In fact, the optimal predictor would often output low-confidence answers. In contrast, prevailing question-answering benchmarks have no data uncertainty, and thus require high confidence for an accurate model to be calibrated. Underspecified, non-realizable prediction tasks allow us to circumvent such potential confounding between calibration and accuracy.
@@ 3.2 Extracting risk scores
To extract risk scores from LLMs, we map each inference problem to a natural text prompt. For a given data point, we specify the classification task in the prompt and extract class probabilities from the model’s next token probabilities, similar to traditional question-answering interfaces. The prompt consists of three components (as shown in Figure 1):
@@
    ” This step is typically not needed in supervised classification tasks, because the training data implicitly defines the population. However, for LLMs this is particularly important for evaluating calibration outside the realizable setting, as risk scores cannot in general simultaneously be calibrated to different populations. Skipping this step could provide insights related to alignment [22, 45, 46] rather than calibration.
@@
When using the constructed multiple-choice prompts, we query the models and extract scores from the next token probabilities for the choice labels A,B as $r_{i}=\mathbb{P}(\texttt{A})/(\mathbb{P}(\texttt{A})+\mathbb{P}(\texttt{B}))$, following the methodology of standard question-answering benchmarks [39, 45]. As LLMs are known to have ordering biases in multiple-choice question-answering [47, 65], we evaluate responses on all choice orderings and average the resulting scores. When using numeric prompting, we prefix the answer with ‘0.’ to improve the likelihood of a direct numeric response, and run two forward passes, selecting the highest likelihood numeric token at each iteration. We refer to Xiong et al. [37] for an overview on alternative design choices on how to elicit confidence scores from LLMs.
@@ 3.3 The folktexts package
The folktexts package is designed to offer a flexible interface between tabular prediction tasks and natural language question-answering tasks in order to extract risk scores from LLMs. The package makes available the ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, and ACSTravelTime prediction tasks [9] as natural language benchmarks, together with various functionalities to customize the task definitions (e.g., use a different set of features to predict income) and subsample the reference data (e.g., predict income only among college graduates in California). The set of attributes available to define the features and label can be found in the ACS PUMS data dictionary.444 https://www.census.gov/programs-surveys/acs/microdata/documentation.html. Additionally, folktexts is compatible with open-source models running locally, as well as with closed-source models hosted through a web API. However, APIs must make available the next token probabilities for each forward pass instead of returning discrete text completions — the OpenAI API is one example that is compatible with folktexts out of the box.
In addition to providing a reproducible way to extract risk scores from LLMs, folktexts also offers pre-implemented evaluation metrics to benchmark and compare the calibration and accuracy of LLMs, as well as easy plotting of group-conditional calibration curves for a cursory view of potential biases. The package is easy to use within a python notebook, as a dependency, or directly from the command line. Further details on usage and design choices are available in Appendix C.